OpenStudy (anonymous):

solve for x 500e^(3x + 2) = 2500

OpenStudy (anonymous):

e^(3x+2)=5 ln(e^(3x+2)=ln5 3x+2=ln5 x=(ln5-2)/3

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

need explanation? feel free to ask!

OpenStudy (anonymous):

i follow

OpenStudy (anonymous):

what you did

OpenStudy (anonymous):

glad i could help bro!

OpenStudy (anonymous):

how did you get the 5 on the right side of the equation

OpenStudy (anonymous):

dividing 2500 by 500

OpenStudy (diyadiya):

Divide both sides by 500 :)

OpenStudy (anonymous):

2500/500= 500

OpenStudy (diyadiya):

No its 5

OpenStudy (anonymous):

hahaha my calc is on drugs

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

what about this log2 x + log2 (x-2) = 3 solve for x

OpenStudy (anonymous):

use brains as calc, usually better

OpenStudy (diyadiya):

and btw people can answer closed questions tooo (you asked that in chat)

OpenStudy (anonymous):

log2 x*(x-2)=3 x^2-2*x=8 x'=4 x''=-2

OpenStudy (anonymous):

no, did not use calc

OpenStudy (anonymous):

just have a question about the 8, how did you get it?

OpenStudy (anonymous):

log definition loga b=c <=> c^a=b just did like 2^(log2 x*(x-2))=2^3

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you do this also 3ln(2x + 3) = 27 solve for x

OpenStudy (anonymous):

ln(2x+3)=9 e^(ln(2x+3))=e^9 2x+3=e^9 x=(e^9-3)/2

OpenStudy (anonymous):

3e^(1-x^2) times ln(x) can you find the equation of the tangent line at x= 1

OpenStudy (anonymous):

sorry, didn't see it! gonna see what i can do now

OpenStudy (anonymous):

didn't understand very well what is asked, post it as a separate question?